# A proof of the Kim-Vu sandwich conjecture

### v1

@@ 6.4 Connection properties of $F\setminus K$
Let $1/n\ll 1/C\ll\varepsilon,\sigma\ll 1$. Let $d,\lambda$ and $\mu$ satisfy $\log^{-3}n\leq\lambda,\mu\ll 1$ and $d\geq C\log n$. Let $m\in\mathbb{N}$ satisfy $\varepsilon dn\leq m\leq n^{1+\sigma}$. Let $\delta=2m/n$ and suppose that $\lambda^{2}\mu\delta\geq C\log n$. Let $p=m(\binom{n}{2}-dn/2)^{-1}$. Let $K\sim G_{d}(n)$, form $E$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F=K+E$. Then, with probability $1-o(n^{-14})$, for each disjoint sets $U,V\subset[n]$ with $|U|,|V|\geq\mu n$, we have $e_{F\setminus K}(U,V)=(1\pm 3\lambda)p|U||V|$.
@@
First note that, for each disjoint $U,V\subset[n]$ with $|U|\geq|V|\geq\mu n$, we have, as $\lambda\mu n\geq\sqrt{n}\geq d$ that
$(1-\lambda)|U||V|\leq|U||V|-d|U|-d|V|\leq e_{K_{n}\setminus K}(U,V)\leq|U||V|.$
@@
$(1-\lambda)p|U||V|\leq\mathbb{E}(e_{F\setminus K}(U,V))\leq p|U||V|.$
As $\lambda^{2}p|U||V|\geq\lambda^{2}\cdot(m/n^{2})\cdot|U|(\mu n)=\lambda^{2}\mu|U|m/n=\lambda^{2}\mu\delta|U|/2=\omega(|U|\log n)$, we have by an application of a Chernoff bound that, with probability $1-\exp(-\omega(|U|\log n))$,
$(1-2\lambda)p|U||V|\leq e_{F\setminus K}(U,V)\leq(1+\lambda)p|U||V|.$
Thus, $e_{F\setminus K}(U,V)\neq(1\pm 3\lambda)p|U||V|$ for some $U,V\subset[n]$ with $|U|\geq|V|\geq\mu n$ with probability at most
$\sum_{m_{1}=\mu n}^{n}\sum_{m_{2}=m_{1}}^{n}\binom{n}{m_{1}}\binom{n}{m_{2}}\exp(-\omega(m_{2}\log n))=o(n^{-14}).\qed$
@@
Take the variables as set up in Theorem 6.1, so that $1/n\ll 1/C\ll 1/C_{0}\ll\varepsilon$, $1/C\ll\mu,\sigma\ll 1$, $d\geq C\log n$, $\varepsilon dn\leq m\leq n^{1+\sigma}$, $\delta=2m/n$, and $p=m(\binom{n}{2}-dn/2)^{-1}$. Furthermore, we have
@@
and $\ell_{0}$ is the smallest integer for which $\delta^{(\ell_{0}-1)}d^{(\ell_{0}-1)}\geq n$. Let $C^{\prime}$ satisfy $1/C\ll C^{\prime}\ll 1$.
Let $K\sim K_{d}(n)$, form $E$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F=K+E$. By Lemma 6.3, in order to prove Theorem 6.1, it is sufficient to show that LABEL:eq:degreesinFK:pseudLABEL:prop:F-Kconnection hold with probability $1-o(n^{-14})$.
As $\delta=pn-d\geq pn-pn/(1+\varepsilon)\geq\varepsilon pn$, we have that
@@
where we have used that $\delta\geq 2\varepsilon d\geq 2\varepsilon C\log n$. Thus, by Lemma 5.3, we have that LABEL:eq:degreesinFK:pseud holds with probability $1-o(n^{-14})$.
@@
By Corollary 6.11 with $\lambda^{\prime}=\lambda/\log n$, as $\lambda^{\prime}\delta d\geq\lambda\delta\geq C\varepsilon\log n$, we have that, with probability $1-o(n^{-14})$, LABEL:eq:fewdoubled:K holds whenever $|U|\leq\lambda n/10^{6}\log n$. Similarly, Corollary 6.11 applied with $\lambda$ shows that, with probability $1-o(n^{-14})$, LABEL:eq:fewdoubled:K holds whenever $|U|\leq\lambda n/10^{6}$. Thus, LABEL:eq:fewdoubled:K holds with probability $1-o(n^{-14})$. Furthermore, by Lemmas 6.12 and 6.13, we have that, with probability $1-o(n^{-14})$, LABEL:eq:fewdoubled:F-K holds with probability $1-o(n^{-14})$, while, by Lemma 6.14, LABEL:prop:new holds with probability $1-o(n^{-14})$.
@@
Thus, by Lemma 6.15 with $\mu=\lambda/10^{8}$, LABEL:prop:F-Kconnection holds with probability $1-o(n^{-14})$. This completes the proof that LABEL:eq:degreesinFK:pseudLABEL:prop:F-Kconnection hold with probability $1-o(n^{-14})$, completing the proof of the theorem. ∎
@@ 7 Upper bounds for switching paths
where we have used that $d\leq n^{\sigma}$ and $\lambda=(\eta/C_{0})^{2}$.
Now, reveal the edges next to $x$ in $E$, and let $Z=B\cap N_{F\setminus K}(x)$. Note that $\mathbb{E}|Z|\leq p\eta n/20\leq\eta\delta/10$. As $\eta\delta\geq C_{0}\delta(\log n/\delta)^{1/8}\geq C_{0}\log n$, by a Chernoff bound, we have that, with probability $1-o(n^{-14})$, $|Z|\leq\eta\delta/5$. By Lemma 5.3, with probability $1-o(n^{-14})$, as $\delta\eta^{2}\geq C_{0}^{2}\delta(\log n/\delta)^{1/4}\geq C_{0}^{2}\log n$, we have that $d_{F\setminus K}(v)\leq(1+\eta/20)\delta$. Assuming these properties, we have both that $|Z|\leq\eta\delta/5$ and that the number of $(F\setminus K,K)$-alternating $x,y$-paths of length $2\ell$ in which the second vertex is not in $Z$ is at most
@@
while $|\mathcal{Q}_{i}|\leq|\mathcal{Q}_{i-1}|\cdot(\eta\delta/5)\cdot d\leq(\eta\delta/5)^{i}d^{i}$.
@@ 8 Concentration for switching paths in the non-critical range
Let $1/n\ll 1$. Let $n^{0.99}\leq d\leq n-1$. Then, for all but at most $e^{-n}|\mathcal{K}_{d}(n)|$ graphs $K\in\mathcal{K}_{d}(n)$, we have the following property.
@@
Let $A\subset[n]$ satisfy $|A|=m$ and let $B=[n]\setminus A$. Enumerate $A=\{a_{1},\ldots,a_{m}\}$ and $B=\{b_{1},\ldots,b_{n-m}\}$. For each $i\in[n-m]$, let $d_{i}=d-d_{G[B]}(b_{i})$. Note that $\mathbb{E}(d_{G[B]}(b_{i}))=q(n-m-1)\geq d/2\geq n^{0.99}/2$ for each $i\in[n-m]$. Thus, by a Chernoff bound and a union bound, with probability at least $1-1/2n^{2}$ we have that, for every $i\in[n-m]$,
@@
As $\mathbb{P}(G=K^{\prime})$ is constant on $K^{\prime}\in\mathcal{K}_{d}(n)$, if $K\sim K_{d}(n)$ then the probability that $K$ does not satisfy LABEL:prop:UVgood is, by Claim 8.4 and Claim 8.5, at most
@@
and thus the lemma also holds when $d<n-n^{0.95}$. ∎
@@ 8.2 Concentration for switching paths for large $d$
By Lemma 8.2, with probability $1-o(n^{-10})$, we have the following property when $d\geq n^{0.99}$.
@@
Fix distinct $x,y\in[n]$ and a set $Z\subset[n]$ with $|Z|\leq 4k$. We will show the property we want holds with probability $1-o(n^{-12-C})$ so that the result follows by a union bound. Let $\mathbf{E}$ be the set of $I\subset[N]$ with $|I|=k$ such that if $H_{I}$ is the graph with vertex set $[n]$ and edge set $\{e_{i}:i\in I\}$, then there is at least one $(H_{I},K)$-alternating $x,y$-path of length $2k$ with no interior vertices in $Z$. For each $I\in\mathbf{E}$, let $w(I)$ be the number of $(H_{I},K)$-alternating $x,y$-paths of length $2k$ with no interior vertices in $Z$. Let $Y=\sum_{I\in\mathbf{E}}w(I)\cdot\prod_{i\in I}\mathbf{1}_{\{e_{i}\in E(F^{\prime})\}}$. Observe that $Y$ is the number of $(F^{\prime},K)$-alternating $x,y$-paths of length $2k$ with no interior vertices in $Z$.
@@
On the other hand, letting $v_{0}=x$ and $v_{2k}=y$, as $\eta d=C_{0}\mu d/\log n\geq C_{0}\mu C$, there are at least $(1-\eta/4)d$ choices of $v_{2k-1}\in N_{K}(v_{2k})\setminus(Z\cup\{v_{0}\})$. Then, as $d<n^{0.99}$, for each $1\leq i\leq k-1$ in turn, there are at least $(d-2k-|Z|)(n-d-1-2k-|Z|)\geq(1-\frac{6k}{d})(1-\frac{6k}{n-1-d})d(n-1-d)\geq(1-\frac{\eta}{10k})d(n-d-1)$ choices for $(v_{2i-1},v_{2i})$ so that $v_{2i-1}v_{2i}\in E(K)$ and $v_{2i-1}$ and $v_{2i}$ are distinct from $v_{0},v_{1},\ldots,v_{2i-2},v_{2k-1}$ and $v_{2k}$, and not in $Z$, and $v_{2i-2}v_{2i-1},v_{2i}v_{2k-1}\notin E(K)$. Therefore, we have
@@
Therefore, we can assume that b) $d\geq n^{0.99}$, so that, in particular, R holds. Let $v_{0}=x$ and $v_{2k}=y$. For each $1\leq i\leq 2k-3$, pick $v_{i}\in[n]\setminus(\{v_{2k},v_{0},v_{1},\ldots,v_{i-1}\}\cup Z)$ such that $v_{i}\in N_{K_{n}\setminus K}(v_{i-1})$ if $i$ is odd and $v_{i}\in N_{K}(v_{i-1})$ is even. Note that the number of ways to do this is at least $\left(d-6k\right)^{k-2}\left(n-d-1-6k\right)^{k-1}=(1\pm\frac{\eta}{10k})^{2k-3}d^{k-2}(n-1-d)^{k-1}=(1\pm\eta/4)d^{k-2}(n-1-d)^{k-1}$, where we have used that $d\geq n^{0.99}$ and $n-1-d\geq\delta\geq n^{\sigma}$. Then, pick distinct $v_{2k-2}$ and $v_{2k-1}$ in $n]\setminus(\{v_{2k},v_{0},v_{1},\ldots,v_{2k-3}\}\cup Z)$ with $v_{2k-2}\in N_{K}(v_{2k-3})$, $v_{2k-1}\in N_{K}(v_{2k})$ and $v_{2k-2}v_{2k-1}\notin E(K)$, so that, by R, the number of choices for $v_{2k-2}$ and $v_{2k-1}$ is, as $d\geq n^{0.99}$ and $n-1-d\geq\delta\geq n^{\sigma}$,
$(1\pm 2n^{-0.01})d^{2}-(1\pm n^{-0.01})d^{2}\cdot\frac{d}{n}=\left(1\pm\frac{\eta}{10}\right)\frac{d^{2}(n-d-1)}{n}.$
@@
Note that, as $\delta\geq\varepsilon d$, and $\delta+d\leq n-1$, we have $d\leq(n-1)/(1+\varepsilon)$, so that $n-1-d\geq\varepsilon d$. Thus, $n-1-d\geq\max\{n-1-d,\varepsilon d\}\geq\varepsilon n/4$. Therefore, including our initial choices, from (51) and (52), we have that

### v2

@@ 6.4 Connection properties of $F\setminus K$
Let $1/n\ll 1/C\ll\varepsilon,\sigma\ll 1$. Let $d,\lambda$ and $\mu$ satisfy $\log^{-3}n\leq\lambda,\mu\ll 1$ and $d\geq C\log n$. Let $m\in\mathbb{N}$ satisfy $\varepsilon dn\leq m\leq n^{1+\sigma}$. Let $\delta=2m/n$ and suppose that $\lambda^{2}\mu\delta\geq C\log n$. Let $p=m(\binom{n}{2}-dn/2)^{-1}$. Let $K\sim G_{d}(n)$, form $E$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F=K+E$. Then, with probability $1-o(n^{-14})$, for each disjoint sets $U,V\subset[n]$ with $|U|,|V|\geq\mu n$, we have $e_{F\setminus K}(U,V)=(1\pm\lambda)\frac{\delta}{n}|U||V|$.
@@
First note that, for each disjoint $U,V\subset[n]$ with $|U|\geq|V|\geq\mu n$, we have, as $\lambda\mu n\geq\sqrt{n}\geq 10d$ that
$\left(1-\frac{\lambda}{5}\right)|U||V|\leq|U||V|-d|U|-d|V|\leq e_{K_{n}\setminus K}(U,V)\leq|U||V|.$
@@
$\left(1-\frac{\lambda}{5}\right)p|U||V|\leq\mathbb{E}(e_{F\setminus K}(U,V))\leq p|U||V|.$
Note that $p=\frac{\delta}{n-d-1}=\left(1\pm\frac{\lambda}{5}\right)\frac{\delta}{n}$. As $\lambda^{2}p|U||V|\geq\lambda^{2}\cdot\left(1-\frac{\lambda}{5}\right)\frac{\delta}{n}\cdot|U|\cdot\mu n\geq\lambda^{2}\mu\delta|U|/2\geq\frac{C}{2}|U|\log{n}$, we have by an application of a Chernoff bound (Lemma 2.6) that, with probability $1-\exp(-\frac{C}{100}|U|\log{n})$,
$\left(1-\frac{\lambda}{2}\right)p|U||V|\leq e_{F\setminus K}(U,V)\leq\left(1+\frac{\lambda}{2}\right)p|U||V|.$
Thus, $e_{F\setminus K}(U,V)\neq(1\pm\lambda)\frac{\delta}{n}|U||V|$ for some $U,V\subset[n]$ with $|U|\geq|V|\geq\mu n$ with probability at most
$\sum_{m_{1}=\mu n}^{n}\sum_{m_{2}=m_{1}}^{n}\binom{n}{m_{1}}\binom{n}{m_{2}}\exp(-\tfrac{C}{100}m_{2}\log{n})=o(n^{-14}).\qed$
@@
Take the variables as set up in Theorem 6.1, so that $1/n\ll 1/C\ll 1/C_{0}\ll\varepsilon$, $1/C\ll\mu,\sigma\ll 1$, $d\geq C\log n$, $\varepsilon dn/2\leq m\leq n^{1+\sigma}$, $\delta=2m/n$, and $p=m(\binom{n}{2}-dn/2)^{-1}$. Furthermore, we have
@@
and $\ell_{0}$ is the smallest integer for which $\delta^{(\ell_{0}-1)}d^{(\ell_{0}-1)}\geq n$. Let $C^{\prime}$ satisfy $1/C\ll 1/C^{\prime}\ll 1$.
Let $K\sim G_{d}(n)$, form $E$ by including each element of $n]^{(2)}\setminus E(K)$ uniformly at random with probability $p$, and let $F=K+E$. By Lemma 6.3, in order to prove Theorem 6.1, it is sufficient to show that LABEL:eq:degreesinFK:pseudLABEL:prop:F-Kconnection hold with probability $1-o(n^{-14})$.
Note that
@@
where we have used that $\delta\geq\varepsilon d\geq\varepsilon C\log n$. Thus, by Lemma 5.3, we have that LABEL:eq:degreesinFK:pseud holds with probability $1-o(n^{-14})$.
@@
By Corollary 6.11 with $\lambda^{\prime}=\lambda/\log n$ and $C^{\prime}$, as $\lambda^{\prime}\delta d\geq\lambda\delta\geq C^{\prime}\log n$, we have that, with probability $1-o(n^{-14})$, LABEL:eq:fewdoubled:K holds whenever $|U|\leq\lambda n/10^{6}d\log n$. Similarly, Corollary 6.11 applied with $\lambda$ and $C^{\prime}$ shows that, with probability $1-o(n^{-14})$, LABEL:eq:fewdoubled:K holds whenever $|U|\leq\lambda n/10^{6}d$. Thus, LABEL:eq:fewdoubled:K holds with probability $1-o(n^{-14})$. Furthermore, by Lemmas 6.12 and 6.13 applied with $C^{\prime}$, we have that, with probability $1-o(n^{-14})$, LABEL:eq:fewdoubled:F-K holds with probability $1-o(n^{-14})$, while, by Lemma 6.14, LABEL:prop:new holds with probability $1-o(n^{-14})$.
@@
Thus, by Lemma 6.15 with $\mu=\lambda/10^{8}$ and $C^{\prime}$, LABEL:prop:F-Kconnection holds with probability $1-o(n^{-14})$. This completes the proof that LABEL:eq:degreesinFK:pseudLABEL:prop:F-Kconnection hold with probability $1-o(n^{-14})$, completing the proof of the theorem. ∎
@@ 7 Upper bounds for switching paths
where we have used that $d\leq n^{\sigma}\leq\eta\sqrt{n}/100$ and $\lambda=(\eta/C_{0})^{2}$.
Now, reveal the edges next to $x$ in $E$, and let $Z=B\cap N_{F\setminus K}(x)$. Note that $\mathbb{E}|Z|\leq p\eta n/20\leq\eta\delta/10$. As $\eta\delta\geq C_{0}\delta(\log n/\delta)^{1/8}\geq C_{0}\log n$, by a Chernoff bound (Lemma 2.6) we have that, with probability $1-o(n^{-14})$, $|Z|\leq\eta\delta/5$. Choose $C^{\prime}$ satisfying $1/c\ll 1/C^{\prime}\ll 1$ so that $\eta\geq(\delta/\log{n})^{3/8}\sqrt{\frac{\log{n}}{\delta}}\geq(C\varepsilon)^{3/8}\sqrt{\frac{\log{n}}{\delta}}\geq 20C^{\prime}\sqrt{\frac{\log{n}}{\delta}}$. Applying Lemma 5.3, with probability $1-o(n^{-14})$ we have that $d_{F\setminus K}(v)\leq(1+\eta/20)\delta$. Assuming these properties, we have both that $|Z|\leq\eta\delta/5$ and that the number of $(F\setminus K,K)$-alternating $x,y$-paths of length $2\ell$ in which the second vertex is not in $Z$ is at most
@@
while $|\mathcal{Q}_{i}|\leq|\mathcal{Q}_{i-1}|\cdot(\eta\delta/5)\cdot d\leq(\eta\delta d/5)^{i}$.
@@ 8 Concentration for switching paths in the non-critical range
Let $1/n\ll 1$. Let $n^{0.99}\leq d<n$. Then, for all but at most $e^{-n}|\mathcal{K}_{d}(n)|$ graphs $K\in\mathcal{K}_{d}(n)$, we have the following property.
@@
Let $A\subset[n]$ satisfy $|A|=m$ and let $B=[n]\setminus A$. Enumerate $A=\{a_{1},\ldots,a_{m}\}$ and $B=\{b_{1},\ldots,b_{n-m}\}$. For each $i\in[n-m]$, let $d_{i}=d-d_{G[B]}(b_{i})$. Note that $\mathbb{E}(d_{G[B]}(b_{i}))=q(n-m-1)\geq d/2\geq n^{0.99}/2$ for each $i\in[n-m]$. Thus, by a Chernoff bound (Lemma 2.6) and a union bound, with probability at least $1-1/2n^{2}$ we have that, for every $i\in[n-m]$,
@@
As $\mathbb{P}(G=K^{\prime})$ is constant on $K^{\prime}\in\mathcal{K}_{d}(n)$, if $K\sim G_{d}(n)$ then the probability that $K$ does not satisfy LABEL:prop:UVgood is, by Claim 8.4 and Claim 8.5, at most
@@
and thus the lemma also holds when $d<n-n^{0.96}$. ∎
@@ 8.2 Concentration for switching paths for large $d$
Note that as $\varepsilon dn/2\leq m\leq\binom{n}{2}-dn/2$, we have $(1+\varepsilon)d\leq n-1$. Thus, $d\leq n/(1+\varepsilon)\leq(1-\varepsilon/2)n$ and by Lemma 8.2, with probability $1-o(n^{-10})$, we have the following property when $d\geq n^{0.99}$.
@@
Fix distinct $x,y\in[n]$ and a set $Z\subset[n]$ with $|Z|\leq 4k$. We will show the property we want holds with probability $1-o(n^{-12-4k})$ so that the result follows by a union bound. Let $\mathbf{E}$ be the set of $I\subset[N]$ with $|I|=k$ such that if $H_{I}$ is the graph with vertex set $[n]$ and edge set $\{e_{i}:i\in I\}$, then there is at least one $(H_{I},K)$-alternating $x,y$-path of length $2k$ with no interior vertices in $Z$. For each $I\in\mathbf{E}$, let $w(I)$ be the number of $(H_{I},K)$-alternating $x,y$-paths of length $2k$ with no interior vertices in $Z$. Let $Y=\sum_{I\in\mathbf{E}}w(I)\cdot\prod_{i\in I}\mathbf{1}_{\{e_{i}\in E(F^{\prime})\}}$. Observe that $Y$ is the number of $(F^{\prime},K)$-alternating $x,y$-paths of length $2k$ with no interior vertices in $Z$.
@@
On the other hand, letting $v_{0}=x$ and $v_{2k}=y$, as $\eta d=C_{0}\mu d/\log n\geq C_{0}\mu C$, there are at least $(1-\eta/4)d$ choices of $v_{2k-1}\in N_{K}(v_{2k})\setminus(Z\cup\{v_{0}\})$. Then, for each $1\leq i\leq k-1$ in turn, choose $(v_{2i-1},v_{2i})$ so that $v_{2i-1}v_{2i}\in E(K)$ and $v_{2i-1}$ and $v_{2i}$ are distinct from $v_{0},v_{1},\ldots,v_{2i-2},v_{2k-1}$ and $v_{2k}$, and not in $Z$, and $v_{2i-2}v_{2i-1},v_{2i}v_{2k-1}\notin E(K)$. For each such $i$, when we go to pick $(v_{2i-1},v_{2i})$, let
$W_{i}=Z\cup\{v_{0},v_{1},\ldots,v_{2i-2}\}\cup\{v_{2k-1},v_{2k}\}\cup N_{K}(v_{2i-2},v_{2k-1}),$
and note that $|W_{i}|\leq 6k+1+2d$. When we choose $(v_{2i-1},v_{2i})$, any $(v_{2i-1},v_{2i})$ with $v_{2i-1}v_{2i}\in E(K-W_{i})$ has the properties we want, and
$2e(K-W_{i})\geq d(n-2|W_{i}|)\geq\left(1-\frac{\eta}{10k}\right)d(n-d-1),$
where we have used that $d<n^{0.99}$ and $1/n\ll\mu,1/k,1/C_{0}$ and $\eta=C_{0}\mu/\log n$. Therefore, we have
@@
Therefore, we can assume that b) $d\geq n^{0.99}$, so that, in particular, R holds. Note, moreover, that, as $d\leq(1-\varepsilon/2)n$, we have $n-d-1\geq\varepsilon n/3$. Let $v_{0}=x$ and $v_{2k}=y$. For each $1\leq i\leq 2k-3$, pick $v_{i}\in[n]\setminus(\{v_{2k},v_{0},v_{1},\ldots,v_{i-1}\}\cup Z)$ such that $v_{i}\in N_{K_{n}\setminus K}(v_{i-1})$ if $i$ is odd and $v_{i}\in N_{K}(v_{i-1})$ is even. Note that the number of ways to do this is at least $\left(d-6k\right)^{k-2}\left(n-d-1-6k\right)^{k-1}=(1\pm\frac{\eta}{10k})^{2k-3}d^{k-2}(n-1-d)^{k-1}=(1\pm\eta/4)d^{k-2}(n-1-d)^{k-1}$, where we have used that $d\geq n^{0.99}$ and $n-1-d\geq\delta\geq n^{\sigma}$. Then, pick distinct $v_{2k-2}$ and $v_{2k-1}$ in $n]\setminus(\{v_{2k},v_{0},v_{1},\ldots,v_{2k-3}\}\cup Z)$ with $v_{2k-2}\in N_{K}(v_{2k-3})$, $v_{2k-1}\in N_{K}(v_{2k})$ and $v_{2k-2}v_{2k-1}\notin E(K)$, so that, by R, the number of choices for $v_{2k-2}$ and $v_{2k-1}$ is, as $d\geq n^{0.99}$ and $n-1-d\geq\varepsilon n/3$,
$(1\pm 2n^{-0.01})d^{2}-(1\pm n^{-0.01})d^{2}\cdot\frac{d}{n}=d^{2}\cdot\left(\frac{n-d-1}{n}\pm 4n^{-0.01}\right)=\left(1\pm\frac{\eta}{10}\right)\frac{d^{2}(n-d-1)}{n}.$
@@
Recall that $n-1-d\geq\varepsilon n/3$. Therefore, including our initial choices, from (52) and (53), we have that